# Fluctuation induced piezomagnetism in local moment altermagnets

### v1

@@ Abstract
It was recently discovered that, depending on their symmetries, collinear antiferromagnets may break spin degeneracy in momentum space, even in absence of spin-orbit coupling. Such systems, dubbed altermagnets, have electronic bands with a spin-momentum texture set mainly by the combined crystal-magnetic symmetry. This discovery motivates the question which novel physical properties derive from altermagnetic order. Here we show that one consequence of altermagnetic order is a fluctuation-driven piezomagnetic response. Using the checkerboard Heisenberg hamiltonian as a prototypical localized moment altermagnet, we determine its fluctuation induced piezomagnetic coefficient considering temperature induced transversal spin fluctuations. We establish in addition that magnetic fluctuations induce an anisotropic thermal spin conductivity.
@@ I Introduction
In contrast to this, we here show that in localized altermagnets transversal spin fluctuations rather have the opposite effect and actually are the drivers of a piezomagnetic response. To be concrete we using a checkerboard Heisenberg hamiltonian as a model d-wave altermagnet and show that while the fluctuation-driven piezomagnetic response is exponentially small at low temperature when the magnetic modes are gapped, it increases with temperature and thus by thermal magnon occupation, reaching a maximum close to the critical temperature. Analytical expressions for this fluctuation induced piezomagnetic response compare well with numerical simulations in which the magnetic system evolves in time according to the stochastic Landau-Lifshitz-Gilbert equations. We also show that the presence of magnetic fluctuations induce a thermal spin conductivity, which is due to the different magnon branches carrying opposite magnetic moment, thus coupling the spin carried by the heat current to the direction of that current.
@@
II Heisenberg checkerboard altermagnet
@@
Here $J>0$, and, in the first sum, $\bm{R}^{\prime}_{\bm{n}}$ counts the nearest neighbors of $\bm{R}_{\bm{n}}$. Amplitudes of the diagonal interactions $|\tilde{J}_{x,y}|<J$ can generally be different, in this way we take into account the deviation from the altermagnetic limit caused by applied mechanical stress. Additionally, we take into account perpendicular easy-axial anisotropy ($K>0$) and interaction with the magnetic field $\bm{B}=B\bm{e}_{z}$. Here $m_{1z}=\bm{m}_{1}(\bm{R}_{\bm{n}})\cdot\bm{e}_{z}$ and $m_{2z}=\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}})\cdot\bm{e}_{z}$ with $\bm{e}_{z}=\bm{e}_{x}\times\bm{e}_{y}$.
@@
$\begin{split}&\omega_{\nu}(\bm{k})=\omega_{0}\left(F_{\bm{k}}\pm\Omega_{\bm{k}}^{-}\right)\pm\gamma B,\\ &F_{\bm{k}}=\sqrt{\left(1+\frac{\kappa}{2}-\Omega_{\bm{k}}^{+}\right)^{2}-\cos^{2}\frac{k_{x}a_{0}}{2}\cos^{2}\frac{k_{y}a_{0}}{2}},\\ &\Omega_{\bm{k}}^{\pm}=\frac{\epsilon_{x}\sin^{2}\frac{k_{x}a_{0}}{2}\pm\epsilon_{y}\sin^{2}\frac{k_{y}a_{0}}{2}}{2},\end{split}$ (2)
@@
Figure 2: (a) – Dispersion relation (2) within the 1st Brillouin zone for the case $\epsilon_{x}=\epsilon_{y}=0.3$, $\kappa=0.04$, and $B=0$. The value of the splitting between branches is shown by the color code at the bottom. (b) – the comparison of the dispersion (2) with the magnon dispersion obtained using the numerical simulations, for detail see Appendix D. (c) – 2D “surfaces” of constant energy $\omega_{\nu}=0.65\omega_{0}$,
An example of the dispersion relation (2) is shown in Fig. 2, which demonstrates the anisotropic (in $k$-space) splitting of the magnon branches typical for the d-wave altermagnets [3], which for the checkerboard AM are rotated with respect to each other by $\pi/2$. Note the very good agreement with the spectra obtained by means of the numerical simulations, for details see Fig. 2(b) and Appendix D.
Figure 3: The fluctuation induced piezomagnetic coupling constant $\eta$ as a function of temperature is shown for $\epsilon_{x}=-\epsilon_{y}=0.1$ and for three different anisotropy values. The inset demonstrates the asymptotic behavior (5) (dashed lines) in the limit of low temperature. Symbols correspond to the results of numerical simulations.
@@ II.1 Fluctuation induced piezomagnetism
where we introduced the magnetization $\mathcal{M}=M/(L_{x}L_{y})$ with $L_{x}L_{y}$ being the altermagnet area, $\bm{q}=a_{0}\bm{k}$ is the dimensionless wave-vector, $\beta^{\prime}=\hbar\omega_{0}/(k_{\textsc{b}}T)$, and we proceed from the summation to integration over the 1st Brillouin zone, assuming the large size of the magnet. Using (2), one can easily show that $\mathcal{M}\to-\mathcal{M}$ under the interchange $\epsilon_{x}\leftrightarrow\epsilon_{y}$. As a consequence, one has $\mathcal{M}=0$ for $\epsilon_{x}=\epsilon_{y}$, i.e. the magnetization vanishes in the AM limit. This property motivates us to write $\mathcal{M}=\eta(\epsilon_{x}-\epsilon_{y})$ introducing the piezomagnetic coupling constant $\eta$. In the limit of low temperatures ($\beta^{\prime}\gg 1$), we estimate integral (4) by means of the Laplace method [18] and obtain
$\eta\approx\frac{g\mu_{\textsc{b}}}{a_{0}^{2}}f(\epsilon_{x},\epsilon_{y})\frac{\Delta^{2}}{\beta^{\prime}}\ln\frac{1}{1-e^{-\beta^{\prime}\Delta}},$ (5)
where $\Delta=\sqrt{\kappa(1+\frac{\kappa}{4})}$ is the gap size in units $\hbar\omega_{0}$. For $\epsilon_{\alpha}\ll 1$, function $f$ can be approximated as follows $f\approx\frac{2}{\pi}\left[1-\frac{3}{2}(\epsilon_{x}+\epsilon_{y})\left(1+\frac{\kappa}{2}\right)\right]$, for details see Appendix B.
@@
The temperature evolution of $\eta$ is shown in Fig. 3. For the limit $T\to 0$, it is exponentially suppressed, which is a typical behavior for a gapped system. Anisotropy strengthens the magnetization stiffness of the sublattices, suppressing thermal occupation of magnons and thus the emergent magnetic moment. This is more clearly shown in the inset of Fig. 3. For the special gapless case ($\kappa=0$), the piezomagnetic coupling constant has $T^{3}$-dependence in the limit of small temperatures for this 2D system, namely $\eta\approx\frac{g\mu_{\textsc{b}}}{a_{0}^{2}}\frac{6}{\pi}\zeta(3)/\beta^{\prime 3}$ with $\zeta(x)$ being Riemann zeta-function.
@@ II.2 Fluctuation induced thermal spin conductivity
is spin capacity per magnon. Based on definition (7) and dispersion relation (2) one can show that $\sigma_{xy}=0$. For the case $\epsilon_{x}=\epsilon_{y}$, the additional symmetry $\sigma_{xx}=-\sigma_{yy}$ takes place and the conductivity tensor obtains the following form
@@
Figure 4: Diagonal elements of the spin conductivity tensor (7) computed for $\epsilon_{x}=\epsilon_{y}=0.1$ and for three different anisotropy values. The inset demonstrates the asymptotic behavior (10) (dashed lines) in the limit of low temperature.
For the case of low temperature $k_{\textsc{b}}T\ll\hbar\omega_{0}$ we obtain $-\sigma_{xx}\approx\sigma_{yy}\approx\sigma$ with
$\sigma\approx\gamma k_{\textsc{b}}\frac{\omega_{0}\tau_{\text{rlx}}}{2\pi}\Delta^{2}(\epsilon_{x}+\epsilon_{y})\ln\frac{1}{1-e^{-\beta^{\prime}\Delta}}.$ (10)
For the exact expressions for the components of $\sigma_{\alpha\beta}$, see Appendix C.

### v2

@@ Abstract
It was recently discovered that, depending on their symmetries, collinear antiferromagnets may break spin degeneracy in momentum space, even in absence of spin-orbit coupling. Such systems, dubbed altermagnets, have electronic bands with a spin-momentum texture set mainly by the combined crystal-magnetic symmetry. This discovery motivates the question which novel physical properties derive from altermagnetic order. Here we show that one consequence of altermagnetic order is a fluctuation-driven piezomagnetic response. Using two Heisenberg models of $d$-wave altermagnets, a checkerboard one and one for rutiles, we determine the fluctuation induced piezomagnetic coefficients considering temperature induced transversal spin fluctuations. We establish in addition that magnetic fluctuations induce an anisotropic thermal spin conductivity.
@@ I Introduction
In contrast to this, we here show that in localized altermagnets transversal spin fluctuations rather have the opposite effect and actually are the drivers of a piezomagnetic response. To be concrete, we use two Heisenberg models of $d$-wave altermagnet and show that while the fluctuation-driven piezomagnetic response is exponentially small at low temperature when the magnetic modes are gapped, it increases with temperature and thus by thermal magnon occupation, reaching a maximum close to the critical temperature. Analytical expressions for this fluctuation induced piezomagnetic response compare well with numerical simulations in which the magnetic system evolves in time according to the stochastic Landau-Lifshitz-Gilbert equations. Since the applied strain breaks the sublattice symmetry for the considered altermagnetic models, our result confirms a recent prediction of the fluctuation-induced ferrimagnetism in sublattice-imbalanced antiferromagnets [11].
We also show that the presence of magnetic fluctuations induces a thermal spin conductivity, which is due to the different magnon branches carrying opposite magnetic moment, thus coupling the spin carried by the heat current to the direction of that current.
@@
II Heisenberg models of $d$-wave altermagnets
@@
Here $J>0$, and, in the first sum, $\bm{R}^{\prime}_{\bm{n}}$ counts the nearest neighbors of $\bm{R}_{\bm{n}}$. Amplitudes of the diagonal interactions $|\tilde{J}_{x,y}|\ll J$ can generally be different, in this way we take into account the deviation from the altermagnetic limit caused by applied mechanical stress. Additionally, we take into account perpendicular easy-axial anisotropy ($K>0$) and interaction with the magnetic field $\bm{B}=B\bm{e}_{z}$. Here $m_{1z}=\bm{m}_{1}(\bm{R}_{\bm{n}})\cdot\bm{e}_{z}$ and $m_{2z}=\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}})\cdot\bm{e}_{z}$ with $\bm{e}_{z}=\bm{e}_{x}\times\bm{e}_{y}$.
@@
$\omega_{\nu}(\bm{k})=\omega_{0}\left(F_{\bm{k}}\pm\Omega_{\bm{k}}^{-}\right)\pm\gamma B,$ (2a)
$F_{\bm{k}}=\sqrt{\left(1+\frac{\kappa}{2}-\Omega_{\bm{k}}^{+}\right)^{2}-\cos^{2}\frac{k_{x}a_{0}}{2}\cos^{2}\frac{k_{y}a_{0}}{2}},$ (2b)
$\Omega_{\bm{k}}^{\pm}=\frac{\epsilon_{x}\sin^{2}\frac{k_{x}a_{0}}{2}\pm\epsilon_{y}\sin^{2}\frac{k_{y}a_{0}}{2}}{2},$ (2c)
@@
Figure 2: (a) – Dispersion relation (2) within the 1st Brillouin zone for the case $\epsilon_{x}=\epsilon_{y}=-0.3$, $\kappa=0.04$, and $B=0$. The value of the splitting between branches is shown by the color code at the bottom. (b) – the comparison of the dispersion (2) with the magnon dispersion obtained using the numerical simulations, for detail see Appendix B. (c) – 2D “surfaces” of constant energy $\omega_{\nu}=0.95\omega_{0}$,
An example of the dispersion relation (2) is shown in Fig. 2, which demonstrates the anisotropic (in $k$-space) splitting of the magnon branches typical for the d-wave altermagnets [3], which for the checkerboard AM are rotated with respect to each other by $\pi/2$. Note the very good agreement with the spectra obtained by means of the numerical simulations, for details see Fig. 2(b) and Appendix B. Here and below we use negative $\tilde{J}_{\alpha}$ to avoid frustration in the system. However, independent of the sign of $\tilde{J}_{\alpha}$, the presented theory is valid under the condition $|\tilde{J}_{\alpha}|\ll J$, which preserves the perpendicular AFM state as the ground state.
Figure 3: Model of a rutile bilayer film. Magnetic atoms belonging to two different layers (and different sublattices $\bm{m}_{1,2}$ in the same time) are shown by blue and yellow balls. The nearest neighbours antiferromagnetic ($J$) and ferromagnetic ($J_{\textsc{fm}}$) Heisenberg exchange couplings act between sublattices and within each of the sublattices, respectively. The altermagnetic properties are modeled by the next-nearest exchange couplings $\tilde{J}_{1,2}$ acting within each of the sublattices. The magnetic unit cell is shown by the green square with side $a_{0}$.
Figure 4: The fluctuation induced piezomagnetic coupling constant $\eta$ as a function of temperature is shown for $\epsilon_{x}=-0.4$, $\epsilon_{y}=-0.2$, and for three different anisotropy values. Symbols correspond to the results of numerical simulations. Dashed lines show the temperature dependence of the normalized ferrimagnetic piezomagnetic constant, $\eta_{\textsc{fm}}/(g\mu_{\textsc{b}}/a_{0}^{2})\propto|\bm{n}|$ for $\mu_{s}/(g\mu_{\textsc{b}})=0.02$, see Appendix E. The temperature dependence of $|\bm{n}|$ is the same as in Fig. 7. Néel temperatures corresponding to different anisotropies are indicated by vertical gray lines. The inset demonstrates the asymptotic behavior (6) (dashed lines) in the limit of low temperature.
As an alternative to the checkerboard, we consider a model of bilayer rutile film shown in Fig. 3. Here we generalize the previously proposed [13] model of rutile assuming that the altermagnetic bonds in different sublattices are generally different $\tilde{J}_{1}\neq\tilde{J}_{2}$. In this manner, we model an effect of strain applied one of the lattice diagonals. Details of the rutile Hamiltonian are explained Appendix C. Applying the same routine as for the checkerboard model Hamiltonian (see Appendix A) we obtain that the dispersion relation of magnons excited over the perpendicular AFM state (shown in Fig. 3) has the same form as Eq. (2a) with
$F_{\bm{k}}=\biggl{\{}\biggl{[}1+\frac{\kappa}{2}+\upsilon\left(\sin^{2}\frac{k_{x}a_{0}}{2}+\sin^{2}\frac{k_{y}a_{0}}{2}\right)$ (3a)
$+\delta\epsilon\sin(k_{x}a_{0})\sin(k_{y}a_{0})\biggr{]}^{2}-\cos^{2}\frac{k_{x}a_{0}}{2}\cos^{2}\frac{k_{y}a_{0}}{2}\biggr{\}}^{1/2}$
$\Omega_{\bm{k}}^{-}=\bar{\epsilon}\sin(k_{x}a_{0})\sin(k_{y}a_{0}).$ (3b)
Here $\upsilon=J_{\textsc{fm}}/J$ and we defined $\delta\epsilon=(\epsilon_{1}-\epsilon_{2})/2$, $\bar{\epsilon}=(\epsilon_{1}+\epsilon_{2})/2$ with $\epsilon_{\alpha}=\tilde{J}_{\alpha}/J$. The definitions of $\kappa$ and $\omega_{0}$ are the same as in Eq. (2a).
@@ II.1 Fluctuation induced piezomagnetism
where we introduced the magnetization $\mathcal{M}=M/(L_{x}L_{y})$ with $L_{x}L_{y}$ being the altermagnet area, $\bm{q}=a_{0}\bm{k}$ is the dimensionless wave-vector, $\beta^{\prime}=\hbar\omega_{0}/(k_{\textsc{b}}T)$, and we proceed from the summation to integration over the 1st Brillouin zone, assuming the large size of the magnet. Expression (5) is universal for both checkerboard and rutile models, but different definitions of $F_{\bm{k}}$ and $\Omega_{\bm{k}}^{-}$ provided in (2) and (3), respectively.
For the checkerboard system, using (2), one can easily show that $\mathcal{M}\to-\mathcal{M}$ under the interchange $\epsilon_{x}\leftrightarrow\epsilon_{y}$. As a consequence, one has $\mathcal{M}=0$ for $\epsilon_{x}=\epsilon_{y}$, i.e. the magnetization vanishes in the AM limit. This property motivates us to write $\mathcal{M}=\eta(\epsilon_{x}-\epsilon_{y})$ introducing a coefficient $\eta$ which for convenience we call the piezomagnetic constant  [^2] Strictly speaking, $\eta$ is a part of the piezomagnetic coefficient $\eta_{\text{piezo}}=\eta(\partial\delta\epsilon/\partial\varepsilon)$, where $\delta\epsilon=\epsilon_{x}-\epsilon_{y}$ and $\varepsilon$ is the applied strain. Consideration of the quantity $(\partial\delta\epsilon/\partial\varepsilon)$ is beyond the scope of this paper.. Utilizing this terminology we assume the linear relation between the applied infinitesimal strain and the change of the exchange integrals, implying the basic mechanism of the monotonic (exponential) decay of the direct exchange between localized electrons with the distance [20]. However, in some materials (e.g. CrI3 [21]) the dependence of some exchange integrals on the inter-atomic distance can be nonmonotonous implying the importance of the nonlinear terms.
@@
The temperature evolution of $\eta$ is shown in Fig. 4. For the limit $T\to 0$, it is exponentially suppressed, which is a typical behavior for a gapped system. Anisotropy strengthens the magnetization stiffness of the sublattices, suppressing thermal occupation of magnons and thus the emergent magnetic moment. This is more clearly shown in the inset of Fig. 4. In the limit of low temperatures ($\beta^{\prime}/\Delta\gg 1$), we estimate integral (5) by means of the Laplace method [22]:
$\eta\approx\frac{g\mu_{\textsc{b}}}{a_{0}^{2}}\frac{2}{\pi}f(\epsilon_{x},\epsilon_{y})\frac{\Delta^{2}}{\beta^{\prime}}\ln\frac{1}{1-e^{-\beta^{\prime}\Delta}},$ (6)
where $\Delta=\sqrt{\kappa(1+\frac{\kappa}{4})}$ is the gap size in units $\hbar\omega_{0}$. Function $f$ is defined in (31a). In the limit of small altermagnetism ($\epsilon_{\alpha}\ll 1$), we approximate $f\approx 1+O(\epsilon_{x}+\epsilon_{y})$. For details see Appendix D.
For the rutile model, the temperature dependence of the piezomagnetic coefficient $\eta=\mathcal{M}/(2\delta\epsilon)$ is qualitatively similar to the dependencies shown in Fig. 4. The low-temperature asymptotic coincides with (6) up to the value of function $f=O(\epsilon_{1}+\epsilon_{2})$, for details see Appendix C. Thus, in the limit of small altermagnetism, one has $\mathcal{M}\propto\tilde{J}_{1}^{2}-\tilde{J}_{2}^{2}$ and $\mathcal{M}\propto\tilde{J}_{x}-\tilde{J}_{y}$ for the rutile and checkerboard models, respectively. And this is the only principle difference between piezomagnetic properties of these models.
As an example of the altermagnet with the symmetry of rutile, we consider RuO2 with the parameters listed in supplemental materials of Ref. 13 and estimate $\eta\approx 0.13\mu_{\textsc{b}}/\mathrm{nm}^{2}$ for the room temperature.
To verify our analytical prediction, we perform spin-lattice simulations for the checkerboard model. The simulations are based on the classical Landau-Lifshitz equations with the thermal fluctuations included. Details of the numerical routine are explained in Appendix B. One should note a nice agreement with analytics in the intermediate range of temperatures, see Fig 4. However, in vicinity of the Néel temperature $T\lessapprox T_{\textsc{n}}$, the simulated piezomagnetic constant drops down reflecting the vanishing of the order parameter $|\bm{n}|$, see Fig. 7. This is in contrast to the analytical result (5) which was derived for the ground state with $|\bm{n}|=1$. Furthermore, we treat magnons as a gas of non-interacting bosons, which is not true at high temperatures. Since the simulations are based on the purely classical dynamics of magnetic moments, the Bose-Einstein statistic for magnons is not incorporated. This gives rise to small discrepancies for small temperatures.
Using the Landau phenomenological theory of altermagnetism [9], one can show that the situation $\tilde{J}_{x}\neq\tilde{J}_{y}$ leads to ferrimagnetism, i.e. $|\bm{m}_{1}|\neq|\bm{m}_{2}|$, with the corresponding piezomagnetic constant $\eta_{\textsc{fm}}=\mu_{s}|\bm{\mathrm{n}}|/(2a_{0}^{2})$, see Appendix E. Using the spin-lattice simulations, we determine the temperature evolution of $|\bm{n}|$ and fit it by the dependence $|\bm{n}|_{\mathrm{fit}}=(1-T/T_{\textsc{n}})^{\alpha}$, see Fig. 7. The temperature dependencies of $\eta_{\textsc{fm}}$ is shown in Fig. 4 by dashed lines. One can see that for the small enough ratio $\mu_{s}/(g\mu_{\textsc{b}})$, the fluctuation-induced piezomagnetism dominates in a wide region of high temperatures.
The gapless limit ($\kappa=0$) is not achievable in the 2D geometry because of the Mermin-Wagner theorem. To consider this limit, we performed a simple 3D generalization of the checkerboard model in which we consider a stacked system composed of the 2D checkerboard layers shown in Fig. 1 with ferromagnetic inter-layer exchange coupling $J_{z}$, see Appendix F. The corresponding low-temperature asymptotic for $\eta^{\text{3D}}$ is presented in Eqs. (40). For the gapless case we estimate $\eta^{\text{3D}}\propto T^{4}$, namely
$\eta^{\text{3D}}\approx\frac{g\mu_{\textsc{b}}}{a_{0}^{2}c_{0}}\frac{8\sqrt{2}\pi^{2}}{45\sqrt{\upsilon_{z}}}\frac{1}{\beta^{\prime 4}}.$ (7)
Here $\upsilon_{z}=J_{z}/J$, $c_{0}$ is the distance between the stacked layers, and we assumed that $k_{\textsc{b}}T\ll\sqrt{\upsilon_{z}}\hbar\omega_{0}$. Note that the $T^{4}$-dependence of the fluctuation-induced magnetization is typical for collinear Heisenberg antiferromagnets with the broken sublattice symmetry [11].
@@ II.2 Fluctuation induced thermal spin conductivity
is spin capacity per magnon. For the dispersion relation (2a), the general form of the thermal spin conductivity tensor defined in (9) is
$\begin{split}&\sigma_{\alpha\beta}=-\frac{\gamma k_{\textsc{b}}}{4}\frac{\tau_{\text{rlx}}\omega_{0}}{L_{x}L_{y}}\beta^{\prime 2}\sum\limits_{\bm{k}\in 1.\text{BZ}}\sum\limits_{\nu=\pm 1}\\ &\nu\frac{(F_{\bm{k}}+\nu\Omega_{\bm{k}}^{-})(\partial_{k_{\alpha}}F_{\bm{k}}+\nu\partial_{k_{\alpha}}\Omega_{\bm{k}}^{-})(\partial_{k_{\beta}}F_{\bm{k}}+\nu\partial_{k_{\beta}}\Omega_{\bm{k}}^{-})}{\sinh^{2}[\frac{\beta^{\prime}}{2}(F_{\bm{k}}+\nu\Omega_{\bm{k}}^{-})]}.\end{split}$ (11)
Expression (11) is universal for both checkerboard and rutile models but definitions of $F_{\bm{k}}$ and $\Omega_{\bm{k}}^{-}$.
For the checkerboard model, $F_{\bm{k}}$, $\Omega_{\bm{k}}^{-}$ are even and $\partial_{k_{\alpha}}F_{\bm{k}}$, $\partial_{k_{\alpha}}\Omega_{\bm{k}}^{-}$ are odd functions of $k_{\alpha}$. From (11) we conclude that the diagonal components $\sigma_{xy}=\sigma_{yx}=0$ are averaging out during summation over $k_{\alpha}$. For the case $\epsilon_{x}=\epsilon_{y}$, additional symmetries $F_{\bm{k}}\to F_{\bm{k}}$ and $\Omega^{-}_{\bm{k}}\to-\Omega^{-}_{\bm{k}}$ take place under the interchange of variables $k_{x}\leftrightarrow k_{y}$. In this case, Eq. (11) results in $\sigma_{xx}=-\sigma_{yy}$ and the conductivity tensor obtains the following form
@@
Figure 5: Diagonal elements of the spin conductivity tensor (9) computed for $\epsilon_{x}=\epsilon_{y}=-0.1$ and for three different anisotropy values. The inset demonstrates the asymptotic behavior (13) (dashed lines) in the limit of low temperature.
For the case of low temperature ($\beta^{\prime}/\Delta\gg 1$) we obtain $-\sigma_{xx}\approx\sigma_{yy}\approx\sigma$ with
$\sigma\approx\gamma k_{\textsc{b}}\frac{\omega_{0}\tau_{\text{rlx}}}{\pi}\Delta^{2}\bar{\epsilon}\ln\frac{1}{1-e^{-\beta^{\prime}\Delta}},$ (13)
where $\bar{\epsilon}=(\epsilon_{x}+\epsilon_{y})/2$. For the exact expressions for the components of $\sigma_{\alpha\beta}$, see Appendix G. For a 3D generalization of the checkerboard model we consider the gapless limit in which $\sigma\propto T^{3}$, namely
$\sigma^{\mathrm{3D}}\approx\gamma k_{\textsc{b}}\frac{\tau_{\text{rlx}}\omega_{0}\epsilon}{c_{0}\sqrt{\upsilon_{z}}\beta^{\prime 3}}\frac{8\sqrt{2}\pi^{2}}{75}.$ (14)
see Appendix F for details. Here we assumed that $\epsilon_{x}=\epsilon_{y}=\epsilon$ and $k_{\textsc{b}}T\ll\sqrt{\upsilon_{z}}\hbar\omega_{0}$.
For the rutile model, tensor $\sigma_{\alpha\beta}$ has the same structure as (12) up the rotation of the reference frame by $\pi/4$, i.e. $\sigma_{xx}=\sigma_{yy}=0$ and $\sigma_{xy}=\sigma_{yx}=\sigma$. The low-temperature behavior of $\sigma$ coincides with (13) up to a numerical constant, see Eq. (30). Since the low-temperature asymptotics of $\eta$ and $\sigma$ obtained for both models coincides up to a constant multiplier, we conjecture that the results (6) and (13) are universal for two-dimansional $d$-wave altermagnets.